# Cosmological Zoom-In Simulations of Milky Way Host Size Dark Matter Halos with a Blue-Tilted Primordial Power Spectrum

### v1

@@ Abstract
Recent observations from the James Webb Space Telescope revealed a surprisingly large number of galaxies formed at high redshift. Along with strong lensing studies and nearby galaxy observations, these could challenge the standard Lambda Cold Dark Matter cosmology with a power-law primordial power spectrum. In this study, we conduct high-resolution cosmological zoom-in dark matter-only simulations of Milky Way host size halos with a blue, tilted primordial power spectrum ($P(k)\propto k^{m_{s}}$ with $m_{s}>1$ at small scales $>1~{}{\rm Mpc}^{-1}$). We find that the blue-tilted subhalo mass functions can be enhanced by more than a factor of two for subhalo masses $M_{\rm sub}\lesssim 10^{10}~{}{\rm M}_{\odot}$, whereas the subhalo $V_{\rm max}$ functions can be enhanced by a factor of four for maximum circular velocities $V_{\rm max}\lesssim 30~{}{\rm km/s}$. The blue-tilted scaled cumulative substructure fraction can be an order of magnitude higher at $\sim$10of the virial radius. The blue-tilted subhalos also have higher central densities, since the blue-tilted subhalos reach the same $V_{\rm max}$ at a smaller distance $R_{\rm max}$ from the center. We have also verified these findings with higher-resolution simulations.
@@ I Introduction
The standard cosmology model includes the single-field slow-roll inflation model Guth (1981); Liddle and Lyth (1993); Steinhardt and Turner (1984); Salopek and Bond (1990); Liddle and Lyth (1992), which predicts the approximately single power-law (PL) primordial power spectrum (PPS) (Lyth and Riotto, 1999), and the lambda cold dark matter model (LCDM, Peebles (1993); Frenk et al. (1985)), which governs the later evolution of our universe. It has achieved great successes on large scales, e.g. supported by the cosmic microwave background (CMB) Planck Collaboration et al. (2020a, b), galaxy surveys Troxel et al. (2018); Blanton et al. (2017) and Lyman-alpha forest Chabanier et al. (2019). However, there are relatively sparse constraints for the scales smaller than $\sim 1~{}{\rm Mpc}^{-1}$. In particular, there have been active debates over the so-called small-scale crisis of LCDM Bullock and Boylan-Kolchin (2017), including the Missing Satellite Problem (MSP, (Klypin et al., 1999; Moore et al., 1999)), the Core-Cuspy problem (Salucci and Burkert, 2000; Flores and Primack, 1994), and the Too-Big-To-Fail problem (Boylan-Kolchin et al., 2011), etc. This leaves room for other models beyond the standard cosmology, including different models of dark matter, effects of baryons, and modifications to the PPS.
@@
One possible modification is to enhance the PPS on small scales (${\rm>1~{}{\rm Mpc}^{-1}}$) (Covi and Lyth, 1999; Martin and Brandenberger, 2001; Gong and Sasaki, 2011), deviating from the PL PPS. In these small-scale enhanced models, PPS will follow the traditional PL model on large scales, preserving the success of the PL PPS, while having an enhancement on small scales. For example, Ref. Hirano et al. (2015) considered blue-tilted (BT) models, which are broken power law PPSs with a higher power law index in scales $>1~{}{\rm Mpc}^{-1}$. They found a significant boost in dark matter halos with BT BBS, especially at high redshift. More complicated forms also exist, such as several bumps in the PPS Tkachev et al. (2024).
With the recently launched James Webb Space Telescope (JWST) Gardner et al. (2006), Ref. Labbé et al. (2023) found a large number of galaxies formed in the high redshift universe ($z>10$), which may be in tension with the standard cosmology Boylan-Kolchin (2023); Lovell et al. (2023). Ref. Parashari and Laha (2023) eased this tension with the BT formalism, based on the Press-Schechter formalism Press and Schechter (1974) and the Sheth-Tormen correction Sheth and Tormen (1999). Ref. Hirano and Yoshida (2024) performed cosmological N-body simulations with the BT model, achieving similar conclusions that the BT model could match the JWST observations without invoking ultra-high star formation rates. However, other mechanisms in galaxy formation models (e.g. Refs. Dekel et al. (2023); Haslbauer et al. (2022); Yung et al. (2024)) could potentially explain the JWST data. Thus, apart from the high redshift massive galaxies, additional observations are required to distinguish the BT and PL models.
@@
The strong lensing observations can provide another constraint. The asymptotic flux ratio relation, which serves as a general principle in smooth lensing potentials (Mao, 1992), is often observed to be violated. This is taken as evidence for the presence of the substructures around lensing galaxies Mao and Schneider (1998). However, in the PL LCDM cosmology, there could be not enough low mass substructures in lensing galaxies to explain the violations Macciò and Miranda (2006); Xu et al. (2009, 2015). This could suggest extra structures not present in dark matter-only simulations under the standard PL LCDM cosmology. On the other hand, Ref. Gilman et al. (2022) constrained sub-galactic PPS to be consistent with PL with the 11 strongly lensed and quadruply image quasars, albeit with possibilities with blue/red tilted PPS on small scales.
The density profiles and statistics of nearby low-mass dwarf galaxies can also inform us about the nature of dark matter, baryonic physics, and the PPS. For example, the observed central densities of dwarf galaxies can provide insights into the nature of dark matter Rocha et al. (2013); Yang et al. (2024) or constraining baryonic physics Di Cintio et al. (2014); Chan et al. (2015); Tollet et al. (2016); Lazar et al. (2020). The “missing satellite problem” describes the discrepancy between the number of subhalos from cosmology dark matter simulations and the observed satellite galaxy population (Klypin et al., 1999; Moore et al., 1999). This problem could signify a challenge to the standard PL LCDM cosmology, or the importance of including accurate galaxy formation physics in dwarf galaxies. Nearby dwarf galaxies also provide constraints on the PPS. For example, Refs. Esteban et al. (2024); Dekker and Kravtsov (2024) employed semi-analytic models to constrain the BT PPS parameter space, using the observed relation between half-light radius and V-band luminosity function in nearby dwarf galaxies.
@@
Thanks to the recent improvements in observations, many more ($\gtrsim 50$) faint satellite galaxies have been discovered in Milky Way (MW) and nearby MW mass hosts Drlica-Wagner et al. (2015); Bechtol et al. (2015); Koposov et al. (2015); Homma et al. (2024); Müller et al. (2024). After correcting for the detection efficiency (completeness correction), Ref. Kim et al. (2018) argued that the “missing satellites problem” can be solved. It could even become the “too many satellites problem”, meaning that there are too many satellite galaxies inferred from observations than expected in simulations/models Kim et al. (2018). This happens if some mechanisms can reduce the number of satellites in simulations /models. For example, a baryonic disk can tidally strip subhalos Garrison-Kimmel et al. (2017), and/or the minimum subhalo mass to host an observable galaxy is large Graus et al. (2019). This “too many satellites problem” motivates us to explore mechanisms to boost the number of subhalos, which includes the BT model.
@@
Motivated by high redshift, strong lensing, and nearby galaxy observations, we explore the impact of the BT PPS on the substructures in dark matter halos, which, to our knowledge, has not been previously investigated with numerical simulations. We perform cosmological zoom-in dark matter-only simulations of MW host size halos with BT PPS to study its substructures at $z=0$. Note that we only modify the PPS but still consider only cold dark matter (i.e. not warm or self-interacting dark matter).
@@
In section II, we will give a theoretical overview, laying down the framework for the BT PPS and our simulation. Then in section III, we introduce our numerical codes for initial conditions, cosmological simulations, and halo identification. We present our main results in section IV and conclude in section V. Additionally, as the first work to use HBT-HERONS (see subsection III.3) for zoom-in simulations, we show the comparison between HBT+HERONS and another halo finder VELOCIraptor in Appendix A. We also run higher-resolution simulations and compare the fiducial results in Appendix B.
@@ II Theoretical Overview
In this paper, along with the PL model, we pick two BT models from Ref. Hirano and Yoshida (2024) [^1] For $k_{p}$, the value in our table is a float number to several digits because original models use $h~{}\rm{Mpc}^{-1}$ as a unit of $k_{p}$, but we convert it into ${\rm Mpc}^{-1}$.: BT_deep and BT_soft, listed in Table 1. BT_soft with a reasonable star formation efficiency $\sim 0.2$ can explain the large number of galaxies formed at high redshift as observed by JWST Hirano and Yoshida (2024). However, to match the JWST observations, BT_deep requires a high star formation efficiency, comparable to PL Hirano and Yoshida (2024). Although BT_deep does not help to explain the JWST observations, it is still interesting to examine its impact on the substructures of dark matter halo.
@@
For all of our PPS models, the matter power spectra at z=1089 and the parameter settings are shown in Figure 2 and Table 1 respectively. BT_soft deviates from PL at $0.7~{}{\rm Mpc}^{-1}$, whereas BT_deep deviates at $3.51~{}{\rm Mpc}^{-1}$, but they have the same power law index afterward.
@@
BT_deep $k_{p}=3.51\ \rm{Mpc}^{-1}$ $m_{s}=1.5$
BT_soft $k_{p}=0.702\ \rm{Mpc}^{-1}$ $m_{s}=1.5$
@@
The wave number $k$ can be connected with a mass scale $M_{l}$ at z=0 Bullock and Boylan-Kolchin (2017), by assuming the sphere radius $r_{l}$ to be half wavelength $\frac{\pi}{k}$:
$M_{l}=$ $\frac{4\pi}{3}{r_{l}}^{3}\rho_{m}=\frac{\Omega_{m}H_{0}^{2}}{2G}r_{l}^{3}$
$=$ $1.71\times 10^{11}\left(\frac{\Omega_{m}}{0.3}\right)\left(\frac{H_{0}}{70}\right)^{2}\left(\frac{r_{l}}{1~{}\rm{Mpc}}\right)^{3}~{}{\rm M}_{\odot}.$ (4)
We further give the pivot mass $M_{p}$ under which the BT model can enhance the structure formation:
$M_{p}=5.29\times 10^{12}\left(\frac{\Omega_{m}}{0.3}\right)\left(\frac{H_{0}}{70}\right)^{2}\left(\frac{k_{p}}{1~{}{\rm Mpc}^{-1}}\right)^{-3}~{}{\rm M}_{\odot}.$ (5)
For the BT_deep and BT_soft models, $M_{p}$ would be $\sim 1.1\times 10^{11}~{}{\rm M}_{\odot}$ and $\sim 1.4\times 10^{13}~{}{\rm M}_{\odot}$ respectively, using the cosmological parameters in our paper (see subsubsection III.2.1). The host halo of an MW-size galaxy is believed to be around $M_{\rm halo}\sim 10^{12}~{}{\rm M}_{\odot}$. Therefore, while both the BT_deep and BT_soft models can enhance the substructures of MW host size halos, only in BT_soft the main halo can be affected.
@@ III.1 Zoom-in initial conditions (IC)
We carry out a suite of cosmological zoom-in simulations of MW host-size halos. We increase the number of particles in a selected zoom-in region, which contains high-resolution (HRS) particles. The particles outside the zoom-in region are called background or low-resolution (LRS) particles. We chose the public code MUSIC - multi-scale cosmological initial conditions (Hahn and Abel, 2011), which can handle multi-scale IC, to generate initial conditions for our cosmological zoom-in simulations. We adopted the Eisenstein and Hu (Eisenstein and Hu, 1998) transfer function and modified it according to section II to mimic the blue-tilted primordial power spectra.
For the convenience of comparison, we adopt the MUSIC configuration files of the existing simulation project FIRE (Hopkins et al., 2018; Wetzel et al., 2023). We choose the zoom-in initial conditions at the size of MW host halos (i.e., $M_{h}\sim 10^{12}~{}{\rm M}_{\odot}$ in mass), specifically, m12i (Wetzel et al., 2016) and m12f (Garrison-Kimmel et al., 2017). We obtain those configuration files from the Flatiron Institute Data Exploration and Comparison Hub  [^3] Publicly available at https://flathub.flatironinstitute.org/fire.
@@ III.2 Simulation code
We perform several zoom-in simulations using the state-of-the-art cosmological simulation code SWIFT (Schaller et al., 2018, 2024) [^4] Publicly available at https://swiftsim.com/. For N-body simulations, SWIFT uses the Fast Multipole Method (FMM) at small scales (Dehnen, 2000), which can reduce the complexity of gravity calculation (including both the potential and force calculation) to $O(N)$. The Particle Mesh Method is also coupled at large scales to handle periodic volumes (Hockney and Eastwood, 1988).

### v2

# Cosmological Zoom-In Simulations of Milky Way Host Mass Dark Matter Halos with a Blue-Tilted Primordial Power Spectrum

@@ Abstract
Recent observations from the James Webb Space Telescope revealed a surprisingly large number of galaxies at high redshift, challenging the standard Lambda Cold Dark Matter cosmology with a power-law primordial power spectrum. Previous studies alleviated this tension with a blue tilted primordial power spectrum ($P(k)\propto k^{m_{s}}$ with $m_{s}>1$ at small scales $>1~{}{\rm cMpc}^{-1}$).
In this study, we examine whether the blue tilted model can boost dark matter substructures especially at low redshift, thereby addressing other potential challenges to the standard cosmology. First, substructures in the standard cosmological model may not be sufficient to explain the anomalous flux ratio problem observed in strong gravitational lensing. Second, the number of observed nearby satellite galaxies could be higher than the theoretical predictions of the standard cosmology, after completeness correction and tidal stripping by baryonic disks.
To study the impact of a blue tilted primordial power spectrum on substructures, we perform high-resolution cosmological zoom-in dark matter-only simulations of Milky Way host mass halos, evolving to redshift $z=0$. At $z=0$, we find that the blue-tilted subhalo mass functions can be enhanced by more than a factor of two for subhalo masses $M_{\rm sub}\lesssim 10^{10}~{}{\rm M}_{\odot}$, whereas the subhalo $V_{\rm max}$ functions can be enhanced by a factor of four for maximum circular velocities $V_{\rm max}\lesssim 30~{}{\rm km/s}$. The blue-tilted scaled cumulative substructure fraction can be an order of magnitude higher at $\sim$10of the virial radius. The blue-tilted subhalos also have higher central densities, since the blue-tilted subhalos reach the same $V_{\rm max}$ at a smaller distance $R_{\rm max}$ from the center. We have also verified these findings with higher-resolution simulations.
@@ I Introduction
The standard cosmology model includes the single-field slow-roll inflation model Guth (1981); Liddle and Lyth (1993); Steinhardt and Turner (1984); Salopek and Bond (1990); Liddle and Lyth (1992), which predicts the approximately single power-law (PL) primordial power spectrum (PPS) (Lyth and Riotto, 1999), and the lambda cold dark matter model (LCDM, Peebles (1993); Frenk et al. (1985)), which governs the later evolution of our universe. It has achieved great successes on large scales, e.g. supported by the cosmic microwave background (CMB) Planck Collaboration et al. (2020a, b), galaxy surveys Troxel et al. (2018); Blanton et al. (2017) and Lyman-alpha forest Chabanier et al. (2019). However, there are relatively sparse constraints for the scales smaller than $\sim 1~{}{\rm cMpc}^{-1}$. In particular, there have been active debates over the so-called small-scale crisis of LCDM Bullock and Boylan-Kolchin (2017), including the Missing Satellite Problem (MSP, (Klypin et al., 1999; Moore et al., 1999)), the Core-Cuspy problem (Salucci and Burkert, 2000; Flores and Primack, 1994), and the Too-Big-To-Fail problem (Boylan-Kolchin et al., 2011), etc. This leaves room for other models beyond the standard cosmology, including different models of dark matter, effects of baryons, and modifications to the PPS.
@@
One possible modification is to enhance the PPS on small scales (${\rm>1~{}{\rm cMpc}^{-1}}$) (Covi and Lyth, 1999; Martin and Brandenberger, 2001; Gong and Sasaki, 2011), deviating from the PL PPS. In these small-scale enhanced models, PPS will follow the traditional PL model on large scales, preserving the success of the PL PPS, while having an enhancement on small scales. For example, Ref. Hirano et al. (2015) considered blue-tilted (BT) models, which are broken power law PPSs with a higher power law index in scales $>1~{}{\rm cMpc}^{-1}$. They found a significant boost in the number of dark matter halos with BT PPS, especially at high redshift. Hence, using the BT model, Ref. Hirano et al. (2015) explained the large number of high redshift massive galaxies observed by the James Webb Space Telescope (JWST) Gardner et al. (2006), which is unexpected from the PL model. More complicated forms of enhancements also exist, such as several bumps in the PPS Tkachev et al. (2024).
Instead of focusing on high redshift, in this study, we examine the impact of the BT model on the low redshift dark matter substructures. There are two major motivations for this: (1) the anomalous flux ratio problem in strong gravitational lensing and (2) potentially the too many satellite problem.
@@
The strong gravitational lensing observations can probe invisible substructures, so it can constrain dark matter substructures and, thus, the cosmological model. For example, the asymptotic flux ratio relation, which serves as a general principle in smooth lensing potentials (Mao, 1992), is often observed to be violated. This is taken as evidence for the presence of the substructures around lensing galaxies (Mao and Schneider, 1998). However, in the PL cosmology, there could be not enough low mass substructures in lensing galaxies to explain the violations (Macciò and Miranda, 2006; Xu et al., 2009, 2015). This could suggest extra structures not present in dark matter-only simulations under the standard PL cosmology. On the other hand, Ref. Gilman et al. (2022) constrained sub-galactic PPS to be consistent with PL with the 11 strongly lensed and quadruply image quasars, albeit with possibilities with blue/red tilted PPS on small scales.
@@
Another motivation for the BT model is that the number of satellite galaxies could be higher than the expected from the PL model. Thanks to the recent improvements in observations, many more ($\gtrsim 50$) faint satellite galaxies have been discovered in Milky Way (MW) and nearby MW mass hosts Drlica-Wagner et al. (2015); Bechtol et al. (2015); Koposov et al. (2015); Homma et al. (2024); Müller et al. (2024). After correcting for the detection efficiency (completeness correction), Ref. Kim et al. (2018) argued that there could be too many satellite galaxies inferred from observations than expected in simulations/models. This happens if some mechanisms can reduce the number of satellite galaxies in simulations / models. For example, a baryonic disk can tidally strip subhalos Garrison-Kimmel et al. (2017), and/or the minimum subhalo mass to host an observable galaxy is large Graus et al. (2019); Jethwa et al. (2018). This “too many satellites problem” motivates us to explore mechanisms to boost the number of subhalos, which includes the BT model.
A related constraint on the primordial power spectrum, though not the focus of this paper, is the density profiles of nearby low-mass dwarf galaxies. For example, the observed central densities of dwarf galaxies can provide insights into the nature of dark matter Rocha et al. (2013); Yang et al. (2024) or constraining baryonic physics Di Cintio et al. (2014); Chan et al. (2015); Tollet et al. (2016); Lazar et al. (2020). Refs. Esteban et al. (2024); Dekker and Kravtsov (2024) employed semi-analytic models to constrain the BT PPS parameter space, using the observed relation between half-light radius and V-band luminosity function in nearby dwarf galaxies.
@@
Motivated by strong lensing, and nearby galaxy observations, we explore the impact of the BT PPS on the substructures in dark matter halos at low redshift, which, to our knowledge, has not been previously investigated with numerical simulations. We perform cosmological zoom-in dark matter-only simulations of MW host mass halos with BT PPS to study its substructures at $z=0$. Note that we only modify the PPS but still consider only cold dark matter (i.e. not warm or self-interacting dark matter).
@@
In section II, we will give a theoretical overview, laying down the framework for the BT PPS and our simulation. Then in section III, we introduce our numerical codes for initial conditions, cosmological simulations, and halo identification. We present our main results in section IV and conclude in section V. Additionally, as the first work to use HBT-HERONS (see subsection III.3) for zoom-in simulations, we show the comparison between HBT-HERONS and another halo finder VELOCIraptor in Appendix A. We also run higher-resolution simulations and compare the fiducial results in Appendix B.
@@ II Theoretical Overview
In this paper, along with the PL model, we pick two BT models from Ref. Hirano and Yoshida (2024) [^1] For $k_{p}$, the value in our table is a float number to several digits because original models use $h~{}\rm{cMpc}^{-1}$ as a unit of $k_{p}$, but we convert it into ${\rm cMpc}^{-1}$.: BT_deep and BT_soft, listed in Table 1. BT_soft with a reasonable star formation efficiency $\sim 0.2$ can explain the large number of galaxies formed at high redshift as observed by JWST Hirano and Yoshida (2024). However, to match the JWST observations, BT_deep requires a high star formation efficiency, comparable to PL Hirano and Yoshida (2024). Although BT_deep does not help to explain the JWST observations, it is still interesting to examine its impact on the substructures of dark matter halo.
@@
For all of our PPS models, the matter power spectra at z=1089 and the parameter settings are shown in Figure 2 and Table 1 respectively. BT_soft deviates from PL at $0.7~{}{\rm cMpc}^{-1}$, whereas BT_deep deviates at $3.51~{}{\rm cMpc}^{-1}$, but they have the same power law index afterward.
@@
BT_deep $k_{p}=3.51\ \rm{cMpc}^{-1}$ $m_{s}=1.5$
BT_soft $k_{p}=0.702\ \rm{cMpc}^{-1}$ $m_{s}=1.5$
@@
We can associate a comoving wavenumber $k$ with a mass scale $M_{l}(z)$ at a given redshift $z$ by considering a sphere with a comoving Lagrangian radius $r_{l}=\pi/k$ Bullock and Boylan-Kolchin (2017):
$M_{l}(z)=$ $\frac{4\pi}{3}{r_{l}}^{3}\rho_{m}(z)$ (4)
$=$ $\frac{4\pi}{3}{r_{l}}^{3}\rho_{crit}(z)\Omega_{m}(z)$
$=$ $\frac{4\pi}{3}{r_{l}}^{3}$
$\rho_{crit,0}\left[\Omega_{m,0}({1+z})^{3}+\Omega_{r,0}({1+z})^{4}+\Omega_{L,0}\right]$
$\frac{\Omega_{m,0}}{\left[\Omega_{m,0}({1+z})^{3}+\Omega_{r,0}({1+z})^{4}+\Omega_{L,0}\right]}$
$=$ $\frac{4\pi}{3}~{}{r_{l}}^{3}~{}\rho_{crit,0}~{}\Omega_{m,0},$ (5)
where $\rho_{crit}(z)$,$\rho_{m}(z)$ are the comoving critical and matter densities respectively. $\Omega_{m}(z)$ is the cosmological density parameter for matter at redshift z. $\Omega_{m,0},\Omega_{r,0},\Omega_{L,0}$ are the cosmological density parameters at $z=0$ for matter, radiation, and dark energy respectively. We assume a flat universe so the curvature density parameter at $z=0$ is zero, i.e. $\Omega_{k,0}=0$. As a result, the mass scale is independent of redshift $z$ because both the comoving Lagrangian radius $r_{l}$ and the comoving average matter density $\rho_{m}(z)$ do not depend on redshift.
We could further name a pivot mass $M_{p}(k_{p})$ for BT model, structures less massive than which could be enhanced while structures more massive should not be affected:
$M_{p}(k_{p})=$ $5.29\times 10^{12}\left(\frac{\Omega_{m,0}}{0.3}\right)\left(\frac{H_{0}}{70}\right)^{2}\left(\frac{k_{p}}{1~{}{\rm cMpc}^{-1}}\right)^{-3}{\rm M}_{\odot}.$ (6)
For the BT_deep and BT_soft models, $M_{p}$ would be $\sim 1.1\times 10^{11}~{}{\rm M}_{\odot}$ and $\sim 1.4\times 10^{13}~{}{\rm M}_{\odot}$ respectively, using the cosmological parameters in our paper (see subsubsection III.2.1). The host halo of an MW-size galaxy is believed to be around $M_{\rm halo}\sim 10^{12}~{}{\rm M}_{\odot}$. Therefore, while both the BT_deep and BT_soft models can enhance the substructures of MW host mass halos, only in BT_soft the main halo can be affected.
@@ III.1 Zoom-in initial conditions (IC)
We carry out a suite of cosmological zoom-in simulations of MW host mass halos. We increase the number of particles in a selected zoom-in region, which contains high-resolution (HRS) particles. The particles outside the zoom-in region are called background or low-resolution (LRS) particles. We chose the public code MUSIC - multi-scale cosmological initial conditions (Hahn and Abel, 2011), which can handle multi-scale IC, to generate initial conditions for our cosmological zoom-in simulations. We adopted the Eisenstein and Hu (Eisenstein and Hu, 1998) transfer function and modified it according to section II to mimic the blue-tilted primordial power spectra.
For the convenience of comparison, we adopt the MUSIC configuration files of the existing simulation project FIRE (Hopkins et al., 2018; Wetzel et al., 2023). We choose the zoom-in initial conditions at the mass of MW host mass halos (i.e., $M_{h}\sim 10^{12}~{}{\rm M}_{\odot}$ in mass), specifically, m12i (Wetzel et al., 2016) and m12f (Garrison-Kimmel et al., 2017). We obtain those configuration files from the Flatiron Institute Data Exploration and Comparison Hub  [^3] Publicly available at https://flathub.flatironinstitute.org/fire.
Several properties of the m12i and m12f host haloes and simulations are listed in Table 2. m12f is 30% more massive and more extended than m12i, resulting in also more substructures. m12i is more concentrated than m12f as expected from its lower halo mass (Diemer and Kravtsov, 2015). m12i has a series of mergers around redshift $z\gtrsim 1$, while m12f is slightly later forming than m12i.
@@ III.2 Simulation code
We perform several zoom-in simulations using the state-of-the-art cosmological simulation code SWIFT (Schaller et al., 2018, 2024)444 publicly available at https://swiftsim.com/. For N-body simulations, SWIFT uses the Fast Multipole Method (FMM) at small scales (Dehnen, 2000), which can reduce the complexity of gravity calculation (including both the potential and force calculation) to $O(N)$. The Particle Mesh Method is also coupled at large scales to handle periodic volumes (Hockney and Eastwood, 1988).